Suppose you are buying your first condo for $145,000, and you will make a $15,000 down payment. You have arranged to finance the remainder $130,000 with a 30-year, monthly payment, amortized mortgage at a 6.5% nominal interest rate, with the first payment due in one month.

What will your monthly payments be?

What will the amount of interest on the first monthly payment be?

What is the amount of principal being paid on the first monthly payment?

What is the balance of principal outstanding at the end of the first monthly payment?
Answer #1
Loan amount borrowed 130000
Annual rate of interest = 6.50%
Monthly rate = 6.50 /12 = 0.5417%
Divide: Annuity PVF at 0.5417% for 360 periiods 158.2042
Monthly instalment 821.72
Interest on first payment = 130000*6.5%*1/12 = 704.17
Principal repayment = 821.72 - 704.17 = 117.55
Principal outstanding at the end of 1st payment = 130000-117.55 = 129882.5